OpenStudy (anonymous):

logbase 2(x+3) + log base 2(x-3)=4 how do i solve for x

OpenStudy (anonymous):

go myininaya :)

myininaya (myininaya):

\[\log_2((x+3)(x-3)=4\] 2^ 2^ \[(x+3)(x-3)=2^4\] \[x^2-9-2^4=0\] \[x^2-9-16=0\] \[x^2-25=0\] x=5,x=-5 but omg no we cant have x=-5 do you know why?

myininaya (myininaya):

because the inside of a log can't be negative or 0 log(-5+3) log(-5-3) both of these insides are negative even it just one was then we would not consider x=-5 as a solution

myininaya (myininaya):

so only solution x=5